OpenStudy (anonymous):

A) Domain, B)Intercepts, C) symmetry, D)Asymptotes, E) Intervals, F) Intervals of Increase or Decrease, G) Concavity and Points of Inflection, H) Sketch the Curve y = sinx/(2 + cosx) A) Domain, B)Intercepts, C) symmetry, D)Asymptotes, E) Intervals, F) Intervals of Increase or Decrease, G) Concavity and Points of Inflection, H) Sketch the Curve y = sinx/(2 + cosx) @Mathematics

OpenStudy (anonymous):

I believe domain is all real numbers

OpenStudy (anonymous):

the numerator domain is all reals, but the denominator cannot equal to zero so when 2 + cos x = 0 then we have division by zero cos x = -2 never occurs though, lol

OpenStudy (anonymous):

ROFL , so the domain is all reals

OpenStudy (anonymous):

intercepts, there are no vertical asymptotes because the denominator is never equal to zero. the numerator is equal to zero when x = pi*k , for k integer

OpenStudy (anonymous):

so x intercepts are pi*k or pi*n

OpenStudy (anonymous):

Intercepts \[y = \frac{\sin x}{\cos {x} + 2} \] \[ \sin x = 0\] \[ x = \sin^{-1}0\] \[y = 0\]

OpenStudy (anonymous):

3) symmetry, it is an odd function. test it by using f(-x) and see what you get out

OpenStudy (anonymous):

ishann, theres more than one x intercept

OpenStudy (anonymous):

what is pi*k or pi*n?

OpenStudy (anonymous):

yea there is @barry

OpenStudy (anonymous):

k stands for an integer, like { ...-3,-2,-1,0,1,2,3,...}

OpenStudy (anonymous):

I only solved it for principle case, General case is barry's solution

OpenStudy (anonymous):

f(-x) = sin(-x) / ( 2 + cos(-x)) = -sin x / ( 2 + cos x ) = - f(x) , so its odd function it has symmetry wrt to origin

OpenStudy (anonymous):

well can you continue solving using principle case Ishaan94

OpenStudy (anonymous):

what about the intervals that are increasing and decreasing, concavity, and points of inflection

OpenStudy (anonymous):

ta, why are you talking to ishaan, i have you the COMPLETE answer ?

OpenStudy (anonymous):

now i wont help

OpenStudy (anonymous):

you can ignore me. i am not helping anymore

OpenStudy (anonymous):

sorry barry2, I was just wondering what was going on, can you please help me?

OpenStudy (anonymous):

ok we are done with symmetry , since you have no questions i can move on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d) asymptotes, there are no vertical asymptotes because the denominator is never equal to zero. so horizontal asymptote?

OpenStudy (anonymous):

0 or is it pi?

OpenStudy (anonymous):

plug in infinity and -infinity

OpenStudy (anonymous):

how am I supposed to plug in infinity?

OpenStudy (anonymous):

you plug it in ( sin oo / ( 2 + cos (oo)

OpenStudy (anonymous):

infinity divided by infinity=infinity

OpenStudy (anonymous):

sin infinity is not infinity

OpenStudy (anonymous):

is it undefine

OpenStudy (anonymous):

well it oscillates between -1 and 1

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

I thought sinx=0 at pi and 0

OpenStudy (anonymous):

we are talking about infinity

OpenStudy (anonymous):

so is -1, 1 the horizontal asymptote? If so why I don't get what you mean by oscillates

OpenStudy (anonymous):

sin x oscillates between 1 and -1 , the graph y = sin x , as x goes to infinity

OpenStudy (anonymous):

what about cos x where does that oscillates?

OpenStudy (anonymous):

also between -1 and 1

OpenStudy (anonymous):

so what will the critical points be to determine the intervals of increasing and decreasing

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok i checked the graph im pretty sure no horizontal asymptote

OpenStudy (anonymous):

I'm about to go to sleep because I got class early in the morning, but I still need help finding the critical points to determine the intervals where in its increasing/decreasing, concavity, and point of inflection

OpenStudy (anonymous):

ok so no asymptotes, we're good

OpenStudy (anonymous):

dont sleep!

OpenStudy (anonymous):

ok, please not to be mean, could you hurry up

OpenStudy (anonymous):

i can , but i have no motivation. wait

OpenStudy (anonymous):

if you paid i work faster. ok , lets see

OpenStudy (anonymous):

y = sinx/(2 + cosx) y ' = [ 2 cos (x) + 1 ] / ( cos x + 2)^2

OpenStudy (anonymous):

y '' = [ 2 sin x ( cos x - 1) ] / ( cos (x) + 2) ^ 3

OpenStudy (anonymous):

what is E) intervals?

OpenStudy (anonymous):

that just talking about finding the critical points just using to test to determine if the graph is increase or decreasing between the points

OpenStudy (anonymous):

ok the critical points occur when y ' = 0 or y ' is undefined

OpenStudy (anonymous):

y ' = 0 when 2 cos x + 1 = 0 , cos x = -1/2, this occurs when

OpenStudy (anonymous):

x = 2pi/3 +2pi*n , 5pi/3 +2pi*n

OpenStudy (anonymous):

those are critical pts

OpenStudy (anonymous):

cos at -pi/3 and -4pi/3 =-1/2

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

cos (-pi/3) = .5

OpenStudy (anonymous):

how does 5pi/3 +2pi*n?

OpenStudy (anonymous):

cos (5pi/3) = -1/2

OpenStudy (anonymous):

and 2pi *n because cos ( x + 2pi*n ) = cos x

OpenStudy (anonymous):

so cos (5pi/3 + 2pi*n) = cos (5pi/3) = -1/2 , same with other solution

OpenStudy (anonymous):

cos(5pi/3) does not equal -1/2, it equal positive 1/2 cos(4pi/3) equal -1/2

OpenStudy (anonymous):

check your calculator, *shrugs*

OpenStudy (anonymous):

cosine is negative in quadrant 2 or 3, 5pi/3 is quadrant 4

OpenStudy (anonymous):

no 5pi/3 is quad 3

OpenStudy (anonymous):

do 5pi/3*180/pi=300

OpenStudy (anonymous):

how much more of this problem is there left?

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

will we still need to determine the intervals of the critcal numbers which I believe is 4pi/3 and 2pi/3

OpenStudy (anonymous):

x = 2pi/3 +2pi*n , 4pi/3 +2pi*n

OpenStudy (anonymous):

I need to go to sleep what are the intervals its increasing and decreasing, and also concavity?

OpenStudy (anonymous):

y = sinx/(2 + cosx) y ' = [ 2 cos (x) + 1 ] / ( cos x + 2)^2 y '' = [ 2 sin x ( cos x - 1) ] / ( cos (x) + 2) ^ 3

OpenStudy (anonymous):

so when is y' > 0 and y' < 0 , 2 cos x + 1 > 0 , cos x > -1/2 ,

OpenStudy (anonymous):

so thats gives a range of -2pi/3< x < 2pi/3 right? for cosx > -1/2

OpenStudy (anonymous):

it is easier to find 2 cos x + 1 < 0 , which is 2pi/3 < x < 4pi/3

OpenStudy (anonymous):

are those the intervals where its increasing/decreasing?

OpenStudy (anonymous):

of course with any multiple of 2pi etc etc.

OpenStudy (anonymous):

so for 2 cos x + 1 > 0 , we have x > 4pi/3 or x < 2pi/3

OpenStudy (anonymous):

right right.

OpenStudy (anonymous):

0 < x < 2pi/3 or when 4pi/3 < x < 2pi it is increasing

OpenStudy (anonymous):

2pi*n < x < 2pi/3 + 2pi (n+1) , or 4pi/3 +2pi*n < x < 2pi + 2pi(n+1)

OpenStudy (anonymous):

decreasing when 2pi/3 +2pi*n< x < 4pi/3+2pi*(n+1)